OpenStudy (anonymous):

Solve the equation. -1/8(x-16) 1/8(x-8)=x+5

OpenStudy (anonymous):

x=3,29

OpenStudy (anonymous):

How do I get to my answer?

OpenStudy (nurali):

\[-\frac{ 1 }{ 8(x-16) }*\frac{ 1 }{ 8(x-8) }=x+5\] Ur question is this?

OpenStudy (anonymous):

Yes.

OpenStudy (nurali):

\[\frac{ -1 }{ 64(x-16)(x-8) }=x+5\] \[\frac{ -1 }{ 64(x^2-8x-16x+128) }=x+5\] \[\frac{ -1 }{ 64(x^2-24x+128) }=x+5\] \[\frac{ -1 }{ 64x^2-1536x+8192 }=x+5\] \[-1=(x+5)(64x^2-1536x+8192)\] \[-1=(64x^3-1536x^2+8192x+320x^2-7680x+40960)\] \[64x^3-1216x^2+512x+40961=0\] Factorize http://www.wolframalpha.com/input/?i=64x%5E3-1216x%5E2%2B512x%2B40961%3D0